For Va=Inf, Find SS voltage gain (Av)
Find input resistance looking into base.

Re=.6kOhm
Rc=5.6kOhm
B=120
Vbeon=.7V
R1=250kOhm
R2=75kOhm


ovu=5v Ai R. Rc ovo use. Skl po Ribi Ioro ai YER
Answer #1

T 5v 5.6kn VR (+) 077 & 0.6kn 7 8=120 R = 250kr Rqz 75kn VB = 5x R2 Rothe - SX 75 25+250 Rea Rill = 250k || isk RB= 250x75 =Small signal Analysis 056 3 { { {Bre Ero Azi 17 - Rill ball prel Rf = R, Il Raa Vi= I pre + I RE Via I Bre+ (8 +1) I RE Zb= vAu= – BRc Bret (B+1) RE Ava – 120x56 79.84 Av=-8.42 Z;= Zoll kill R = 79.84 || 250 |175 - 19.64 || 57.7 2;= 33.5 kr